# All loop soft photon theorems and higher spin currents on the celestial sphere

### v2

@@ III Soft theorems in conformal primary basis
In order to interpret the soft theorems (5) as Ward identities in a CFT2 living on the celestial sphere we transform to the conformal primary basis. We parametrize the soft momentum $k^{\mu}$, positive helicity photon polarization vector $\varepsilon^{+\mu}(k)$ and the momentum $p^{\mu}$ of a massless hard particle as
$\displaystyle\begin{gathered}k^{\mu}=\eta\omega\,\mathbf{n}^{\mu}(w,\bar{w}),\ \omega\rightarrow 0+\\ \mathbf{n}(w,\bar{w})=\big(1+w\bar{w},w+\bar{w},-i(w-\bar{w}),1-w\bar{w}\big)\\ \quad\varepsilon^{+\mu}(k)=\frac{1}{\sqrt{2}}(\bar{w},1,-i,-\bar{w})\\ p^{\mu}=\eta\omega\mathbf{n}(z,\bar{z})\end{gathered}$ (11)
where $\eta=\pm 1$ for an incoming (outgoing) particle and $(w,\bar{w})((z,\bar{z}))$ represent the (stereographic) coordinates of the Celestial sphere. [^6] The charges $\{e_{a}\}$ are related to their actual charge values $\{q_{a}\}$ by relation $e_{a}=\eta_{a}q_{a}$. We also define celestial primary field $\phi^{(\eta)}_{h,\bar{h}}(z,\bar{z},e)$ by Mellin transforming the creation and annihilation operators $b^{\dagger}$ and $b$ as Pasterski:2016qvg
@@
Now after transforming to the conformal primary basis, the soft photon theorems in (5) can be expressed as the following celestial correlation functions:
@@
for $j=1/2,1,3/2,\cdots$. In the above equations we have kept the helicity and all other internal quantum numbers of hard particles implicit because the soft factors depend only on the electric charges and the momenta of them. The soft operator $\mathcal{S}^{\Delta}(w,\bar{w})$ for positive helicity outgoing soft photon has been defined via the Mellin transformation of the photon annihilation operator $a(k,\sigma=+1)$ along with the conformal soft limit Donnay:2018neh taken inside the celestial correlation functions as [^7] Here we used the following result of Mellin transformation $\int_{0}^{\infty}d\omega\,\omega^{s-1}\omega^{n}(\ln\omega)^{k}=\frac{(-1)^{k}k!}{(s+n)^{k+1}}$, as derived in Krishna:2023ukw . One may be tempted to resum all the multi-order poles in the scaling dimensions and interpret the resummed Mellin-space Ward identity as a simple pole in $\Delta$ with anomalous dimensions. However, the resulting structure of the sum appears to be of the form $\sum_{n=0}^{\infty}(\Delta+n)^{n+2}$ in Mellin space, or equivalently exhibits the integral behavior $\int_{0}^{\infty}d\omega\omega^{\omega+\Delta-2}$, which falls into the category of the “Sophomore’s Dream”, for which no closed analytical result is known.
@@
The Weinberg soft theorem (15) is known to be equivalent to the Ward-identity of a $U(1)$ Kac-Moody algebra at level zero. In the rest of the paper we focus on the symmetry interpretation of the all loop soft theorems (16).
@@ IV $\mathcal{O}(\ln\omega)$ soft theorem as Ward identity and the dipole current operator
$\left\langle\mathcal{S}^{0}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}{e_{a}}^{2}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle$ (20)
@@
$\mathcal{S}^{0}(w,\bar{w})\phi^{(\eta_{a})}_{\Delta_{a}}(z_{a},\bar{z}_{a},e_{a})\sim\frac{\eta_{a}e_{a}^{2}}{w-z_{a}}\left(\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\right)\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})$ (21)
@@
$\boxed{\left\langle\bar{d}(\bar{w},\bar{z})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a}e_{a}\frac{\bar{w}-\bar{z}_{a}}{\bar{z}-\bar{z}_{a}}\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle}$ (22)
@@ V Conformal transformation law of the dipole operator
and for infinitesimal translation, dilatation and special conformal transformations (SCT)888 $\displaystyle\text{Translation:}\quad\begin{pmatrix}1&b\\ 0&1\end{pmatrix},\qquad\text{Dilatation:}\quad\begin{pmatrix}\lambda^{\frac{1}{2}}&0\\ 0&\lambda^{-\frac{1}{2}}\end{pmatrix},\qquad\text{SCT:}\quad\begin{pmatrix}1&0\\ -c&1\end{pmatrix}$, (31) we get the following commutation relations
@@ VII Correlation functions of dipole currents and matter fields
where $k$ is a real number which can be zero also. However, as we will see we can take $k$ to be nonzero without changing the single soft photon theorems. Our current understanding of the role of dipole symmetry in the loop level soft theorems does not allow us to determine the value of $k$. This is discussed in more detail in section XI.
@@
VIII $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ soft theorem as Ward identity
In this section, we focus on the singular parts of the soft theorems in (16) which contain poles in $w$. The Weinberg (15) and the singular parts of the logarithmic (16) soft theorems can be combined as
$\displaystyle\begin{gathered}\left\langle\mathcal{S}^{1-2j}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}\\ =\frac{1}{(2j)!}\sum_{a=1}^{N}\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]^{2j}\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2j}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,\end{gathered}$ (63)
where $j=0,\frac{1}{2},1,\cdots$. Note that for $j=\frac{1}{2}$ the singular part coincides with the actual soft theorem whose Ward identity interpretation has been discussed in section IV. In this section we start with the analysis of the (singular part of the) soft theorem for $j=1$ using the notation introduced in section VI.
@@ VIII $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ soft theorem as Ward identity
$\mathcal{S}^{-1}(w,\bar{w})=\frac{1}{2}\bar{W}^{\alpha}\bar{W}^{\beta}\mathcal{S}^{-1}_{\alpha\beta}(w).$ (64)
and rewrite the singular part of the $j=1$ soft theorem as
$\begin{gathered}\left\langle\mathcal{S}^{-1}_{\alpha\beta}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}\\ =\sum_{a=1}^{N}\frac{\eta_{a}^{2}e_{a}^{3}}{w-z_{a}}\left(\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{Z}_{b\alpha}}{\bar{z}_{a}-\bar{z}_{b}}\right)\left(\sum_{\begin{subarray}{c}c=1\\ c\neq a\end{subarray}}^{N}e_{c}\,\frac{\bar{Z}_{c\beta}}{\bar{z}_{a}-\bar{z}_{c}}\right)\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.\end{gathered}$ (65)
Now to interpret (65) as (singular part of) a Ward identity we, by comparing with (57), introduce a spin-1 current $\bar{d}^{1}_{\alpha\beta}(\bar{z})$ such that
$\left\langle\mathcal{S}^{-1}_{\alpha\beta}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}=\sum_{a=1}^{N}\frac{\eta_{a}^{2}e_{a}^{3}}{w-z_{a}}\left\langle:\bar{d}^{1}_{\alpha\beta}\phi^{(\eta_{a})}_{\Delta_{a}-2}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,$ (66)
@@
and using this we can rewrite the singular part (63) of the soft theorem as
$\begin{gathered}\left\langle\mathcal{S}^{1-2j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{\textbf{singular}}\\ =\sum_{a=1}^{N}\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}\left\langle:\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}\phi^{(\eta_{a})}_{\Delta_{a}-2j}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle,\end{gathered}$ (71)
where the definition of the higher spin current $\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(\bar{z})$ is given in (58). The current $\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}$ is completely symmetric in the spinor indices and transforms in the tensor product of two $SL(2,\mathbb{R})_{\text{R}}$ representations one of which is a conformal highest weight representation of weight $\bar{h}=j$ and the other one is the $(2j+1)$-dimensional spin-$j$ representation. We discuss this in more detail in section IX.1.
In the following section we discuss the algebra of these currents for general values of $k$.

### v3

@@ III Soft theorems in conformal primary basis
In order to interpret the soft theorems (8) as Ward identities in a CFT2 living on the celestial sphere we transform to the conformal primary basis. Since we are working in the high energy limit we can treat the hard particles as (effectively) massless because the finite mass correction is suppressed by powers of $m/E$ and hence small according to our assumption. This allows us to parametrize the soft momentum $k^{\mu}$, positive helicity photon polarization vector $\varepsilon^{+\mu}(k)$ and the momentum $p^{\mu}$ of a hard particle as
$\displaystyle\begin{gathered}k^{\mu}=\eta\omega\,\mathbf{n}^{\mu}(w,\bar{w}),\ \omega\rightarrow 0+\,,\\ \mathbf{n}(w,\bar{w})=\big(1+w\bar{w},w+\bar{w},-i(w-\bar{w}),1-w\bar{w}\big),\\ \quad\varepsilon^{+\mu}(k)=\frac{1}{\sqrt{2}}(\bar{w},1,-i,-\bar{w}),\\ p^{\mu}=\eta\omega\mathbf{n}(z,\bar{z}),\end{gathered}$ (14)
where $\eta=\pm 1$ for an incoming (outgoing) particle and $(w,\bar{w})((z,\bar{z}))$ represent the (stereographic) coordinates of the Celestial sphere. We also define celestial primary field $\phi^{(\eta)}_{h,\bar{h}}(z,\bar{z},e)$ by Mellin transforming the creation and annihilation operators $b^{\dagger}$ and $b$ as Pasterski:2016qvg
@@
Now after transforming to the conformal primary basis, the soft photon theorems in (8) can be expressed as the following identities between celestial correlation functions:
@@
for $j=1/2,1,3/2,\cdots$. In the above equations we have kept the helicity and all other internal quantum numbers of hard particles implicit because the soft factors depend only on the electric charges and the momenta of them. The soft operator $\mathcal{S}^{\Delta}(w,\bar{w})$ for positive helicity outgoing soft photon has been defined via the Mellin transformation of the photon annihilation operator $a(k,\sigma=+1)$ along with the conformal soft limit Donnay:2018neh taken inside the celestial correlation functions as [^8] Here we used the following result of Mellin transformation $\int_{0}^{\infty}d\omega\,\omega^{s-1}\omega^{n}(\ln\omega)^{k}=\frac{(-1)^{k}k!}{(s+n)^{k+1}}$, as derived in Krishna:2023ukw ; Guevara:2019ypd . One may be tempted to resum all the multi-order poles in the scaling dimensions and interpret the resummed Mellin-space Ward identity as a simple pole in $\Delta$ with anomalous dimensions. However, the resulting structure of the sum appears to be of the form $\sum_{n=0}^{\infty}(\Delta+n)^{n+2}$ in Mellin space, or equivalently exhibits the integral behavior $\int_{0}^{\infty}d\omega\omega^{\omega+\Delta-2}$, which falls into the category of the “Sophomore’s Dream”, for which no closed analytical result is known.
@@
The Weinberg soft theorem (18) is known Strominger:2017zoo ; Strominger:2013lka to be equivalent to the Ward-identity of a $U(1)$ Kac-Moody algebra at level zero. In the rest of the paper we focus on the symmetry interpretation of the all loop soft theorems (19).
@@ IV $\mathcal{O}(\ln\omega)$ soft theorem as Ward identity and the dipole current operator
$\left\langle\mathcal{S}^{0}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a=1}^{N}\frac{\eta_{a}{e_{a}}^{2}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.$ (23)
@@
$\mathcal{S}^{0}(w,\bar{w})\phi^{(\eta_{a})}_{\Delta_{a}}(z_{a},\bar{z}_{a},e_{a})\sim\frac{\eta_{a}e_{a}^{2}}{w-z_{a}}\left(\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\right)\phi^{(\eta_{a})}_{\Delta_{a}-1}(z_{a},\bar{z}_{a},e_{a}),$ (24)
@@
$\boxed{\left\langle\bar{d}(\bar{w},\bar{z})\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle=\sum_{a}e_{a}\frac{\bar{w}-\bar{z}_{a}}{\bar{z}-\bar{z}_{a}}\left\langle\prod_{i}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle.}$ (25)
@@ V Conformal transformation law of the dipole operator
and for infinitesimal translation, dilatation and special conformal transformations (SCT)999 $\displaystyle\text{Translation:}\quad\begin{pmatrix}1&b\\ 0&1\end{pmatrix},\qquad\text{Dilatation:}\quad\begin{pmatrix}\lambda^{\frac{1}{2}}&0\\ 0&\lambda^{-\frac{1}{2}}\end{pmatrix},\qquad\text{SCT:}\quad\begin{pmatrix}1&0\\ -c&1\end{pmatrix}$, (34) we get the following commutation relations
@@ VII Correlation functions of dipole currents and matter fields
where $k$ is a real number which can be zero also. However, as we will see we can take $k$ to be nonzero without changing the single soft photon theorems. Our current understanding of the role of dipole symmetry in the loop level soft theorems does not allow us to determine the value of $k$. This is discussed in more detail in section X.
@@
VIII $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ soft theorem and Celestial OPE
In this section, we show how to write down the OPE between the soft photon and a hard charged operator using the soft theorems in (19). As $w\rightarrow z_{a}$ the Weinberg (18) and the singular parts of the logarithmic (19) soft theorems can be combined to give
$\displaystyle\begin{gathered}\left\langle\mathcal{S}^{1-2j}(w,\bar{w})\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{w\rightarrow z_{a}}\\ =\frac{1}{(2j)!}\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}\Bigg[\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{w}-\bar{z}_{b}}{\bar{z}_{a}-\bar{z}_{b}}\Bigg]^{2j}\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2j}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle+\text{non-singular terms}\end{gathered}$ (66)
where $j=0,\frac{1}{2},1,\cdots$. Let us first discuss the $j=1$ case.
@@ VIII $\mathcal{O}(\omega^{2j-1}(\ln\omega)^{2j})$ soft theorem and Celestial OPE
$\mathcal{S}^{-1}(w,\bar{w})=\frac{1}{2}\bar{W}^{\alpha}\bar{W}^{\beta}\mathcal{S}^{-1}_{\alpha\beta}(w),$ (67)
and rewrite (66) as
$\begin{gathered}\left\langle\mathcal{S}^{-1}_{\alpha\beta}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{w\rightarrow z_{a}}\\ =\frac{\eta_{a}^{2}e_{a}^{3}}{w-z_{a}}\left(\sum_{\begin{subarray}{c}b=1\\ b\neq a\end{subarray}}^{N}e_{b}\,\frac{\bar{Z}_{b\alpha}}{\bar{z}_{a}-\bar{z}_{b}}\right)\left(\sum_{\begin{subarray}{c}c=1\\ c\neq a\end{subarray}}^{N}e_{c}\,\frac{\bar{Z}_{c\beta}}{\bar{z}_{a}-\bar{z}_{c}}\right)\left\langle\phi^{(\eta_{a})}_{\Delta_{a}-2}(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle\\ +\text{non-singular terms}\end{gathered}$ (68)
Now by comparing with (60), we write (68) as
$\begin{gathered}\left\langle\mathcal{S}^{-1}_{\alpha\beta}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{w\rightarrow z_{a}}=\frac{\eta_{a}^{2}e_{a}^{3}}{w-z_{a}}\left\langle:\bar{d}^{1}_{\alpha\beta}\phi^{(\eta_{a})}_{\Delta_{a}-2}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle\\ +\text{non-singular terms}\end{gathered}$ (69)
@@
and (66) can be rewritten as
$\begin{gathered}\left\langle\mathcal{S}^{1-2j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(w)\prod_{i=1}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle_{w\rightarrow z_{a}}\\ =\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}\left\langle:\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}\phi^{(\eta_{a})}_{\Delta_{a}-2j}:(z_{a},\bar{z}_{a},e_{a})\prod_{\begin{subarray}{c}i=1\\ i\neq a\end{subarray}}^{N}\phi^{(\eta_{i})}_{\Delta_{i}}(z_{i},\bar{z}_{i},e_{i})\right\rangle+\text{non-singular terms}\end{gathered}$ (74)
where the definition of the higher spin current $\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(\bar{z})$ is given in (61). The equation (66) leads to the following OPE between the soft operator $\mathcal{S}^{1-2j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(w)$ and a hard operator $\phi^{(\eta_{a})}_{\Delta_{a}}(z_{a},\bar{z}_{a},e_{a})$
$\mathcal{S}^{1-2j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}(w)\phi^{(\eta_{a})}_{\Delta_{a}}(z_{a},\bar{z}_{a},e_{a})=\frac{\eta_{a}^{2j}e_{a}^{2j+1}}{w-z_{a}}:\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}\phi^{(\eta_{a})}_{\Delta_{a}-2j}:(z_{a},\bar{z}_{a},e_{a})+\text{non-singular terms}$ (75)
The current $\bar{d}^{j}_{\alpha_{1}\alpha_{2}\cdots\alpha_{2j}}$ is completely symmetric in the spinor indices and transforms in the tensor product of two $SL(2,\mathbb{R})_{\text{R}}$ representations one of which is a conformal highest weight representation of weight $\bar{h}=j$ and the other one is the $(2j+1)$-dimensional spin-$j$ representation. We discuss this in more detail in section IX.1.
Now unlike the $j=0$ and $j=\frac{1}{2}$ cases, the higher $j$ soft theorems cannot be written as Ward identities using the higher spin currents alone. However, as equation (75) shows the singular term in the OPE is completely determined in terms of the higher spin current. To write the soft theorems as Ward identities for $j\geq 1$ we need to understand the part of the soft theorem which does not have a pole in $w$. We leave this interesting question for future investigation.
In the following section we discuss the algebra of the higher spin currents for general values of $k$.